OpenStudy (anonymous):

find the term x^4 in the expansion ((x/2)-3)^7

OpenStudy (aum):

\[\large \left(\frac x2 - 3\right)^7 = \sum_{k = 0}^{7}\binom{7}{k}\left( \frac x2\right)^{7-k}(-3)^k \normalsize ~~~~ \text{-------(1)} \\ \text{ } \\ x^4 \text{ occurs when } k = 3. \\ \text{ } \\ \text{Put k = 3 in (1) and evaluate: } \binom{7}{k}\left( \frac x2\right)^{7-k}(-3)^k \]